OpenStudy (anonymous):

Graph the line containing the given pairof points and find the slope (3,0),(-6,-3) I know m= 1/3

OpenStudy (anmolsingh):

dude now use the equation y=mx+b where m is the slope and b is the y-intercept. Now take any one coordinate and substitute the value in the equation shown above and you will find b(the y-intercept). Now start your graph from there. now since m=1/3=rise/run, go 4 units right(positive) and one unit up(positive). that will give you your graph.